A piece of metal, mass 3.6 kg is suspended from a spring balance. What is the reading of the balance (in N) a.      With the metal in air [36N]

b.     With the metal in water [32N]

c.      With the metal in brine [31.2N]


d.     What is the density of a liquid in which it gives a reading of 33N? [750kg/m­­­­­3].

(Density of brine 1200 kg/m3; density of metal=900kg/m3)


1
Expert's answer
2021-03-21T11:27:35-0400

(a)

"N=mg=3.6\ kg\cdot10\ \dfrac{m}{s^2}=36\ N."

(b)

"N=mg-F_B=\rho_mV_mg-\rho_wV_wg,""N=(\rho_mV_m-\rho_wV_w)g=\rho_mV_mg(1-\dfrac{\rho_w}{\rho_m}),""N=mg(1-\dfrac{\rho_w}{\rho_m}),""N=3.6\ kg\cdot10\ \dfrac{m}{s^2}\cdot(1-\dfrac{1000\ \dfrac{kg}{m^3}}{9000\ \dfrac{kg}{m^3}})=32\ N."

(c)

"N=mg(1-\dfrac{\rho_w}{\rho_b}),""N=3.6\ kg\cdot10\ \dfrac{m}{s^2}\cdot(1-\dfrac{1200\ \dfrac{kg}{m^3}}{9000\ \dfrac{kg}{m^3}})=31.2\ N."

(d)

"N=mg(1-\dfrac{\rho_L}{\rho_m}),""\rho_L=(1-\dfrac{N}{mg})\rho_m,""\rho_L=(1-\dfrac{33\ N}{3.6\ kg\cdot10\ \dfrac{m}{s^2}})\cdot9000\ \dfrac{kg}{m^3}=750\ \dfrac{kg}{m^3}."